OpenStudy (anonymous):

Multiply: (3x2)(6x4).

OpenStudy (anonymous):

its squared

OpenStudy (kalynne):

it would be 144

OpenStudy (wwhitlock):

the whole thing is squared? \[((3x2)(6x4))^2\]

OpenStudy (anonymous):

A. 9x8 B. 18x8 C. 18x6 D. None of the above

OpenStudy (wwhitlock):

Use the Equation button (below, left) to make them look right

OpenStudy (anonymous):

those are the choices

OpenStudy (wwhitlock):

It's not squared then. (3x2)(6x4)= (6)(24)= 6x24= 144 that answer is not listed so.....

OpenStudy (shishi990):

b bc 18*8=144 and (3*2)(6*4)=144